X-7y=56+y
0.09y=0.03x-0.78
Solve the system
I don't understand? please anyone help. thanks

Answer by scott8148(6628)

solving the system means finding values for x and y that work in both equations

x-7y=56+y ... x=56+8y ... multiplying the second equation by 100 gives 9y=3x-78

substituting gives 9y=3(56+8y)-78 ... 9y=168+24y-78 ... -15y=90 ... y=-6

x=56+8(-6) ... x=56-48 ... x=8